Diffusion Morning Report 04/10/06

Ted Berg tberg at snf.stanford.edu
Mon Apr 10 06:32:10 PDT 2006


Good Morning All,
    Looking Ok this morning. Nitride not pumping down-looking into it. 
Sample dropped in SEM hitachi-will retreive. The temp
cal on POCl3 should be done today. Everythingelse looks OK for now.


	Date: 	4/10/06 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	
	
	
Tylan 1 	Up 	  	
	
	
Tylan 2 	Up 	  	
	
	
Tylan 3 	Up 	  	
	
	
Tylan 4 	Up 	  	
	
	
Tylan 5-BBr3 	Up 	  	
	
	
Tylan 6-POCl3 	Down 	temp cal 	
	
	
Tylan 7-TEOS2 	Up 	  	16.69/30 	
	
Tylan 8-SiGe 	Up 	  	27.5/30 	
	
Tylan 9-Poly 	Up 	  	19.61/30 	
	
Tylan 10-Nitride 	Up 	  	18.54/30 	
	
Tylan 11-BPSG-LTO 	Up 	  	21.05/30 	
	
Tylan 12-FGA 	Up 	  	  	
	
  	  	
	
	
	
Thermco 1 	Up 	  	
	
	
Thermco 2 	Up 	  	
	
	
Thermco Poly 	Up 	  	
	
	
Thermco Nitride 	Down 	Testing 	
	
	

	
	
	
	
	
Epi 	up 	  	
	
	

	
	
	
	
	
Burn Box 	Up 	  	
	
	
ZYGO 	up 	  	   	
	

	
	
	
	
	
GAAS 22 	UP 	
	
	
	
GAAS 23 	up 	  	
	
	

	
	
	
	
	
SiH4 for furnaces 	
	A on line 645 psi 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060410/b50454c8/attachment.html>


More information about the diffusion mailing list